### v1

# Teaching Data Science Students to Sketch Privacy Designs through Heuristics (Extended Technical Report)

Abstract
Recent studies reveal that experienced data practitioners often draw sketches to facilitate communication around privacy design concepts. However, there is limited understanding of how we can help novice students develop such communication skills. This paper studies methods for lowering novice data science students’ barriers to creating high-quality privacy sketches. We first conducted a need-finding study (N=12) to identify barriers students face when sketching privacy designs. We then used a human-centered design approach to guide the method development, culminating in three simple, text-based heuristics. Our user studies with 24 data science students revealed that simply presenting three heuristics to the participants at the beginning of the study can enhance the coverage of privacy-related design decisions in sketches, reduce the mental effort required for creating sketches, and improve the readability of the final sketches.
00footnotetext:  00footnotetext: Equal contributions.00footnotetext: This report is an extended version of a IEEE S&P 2025 paper. We include Table XIV to further prove the improved communication efficiency of heuristic-guided privacy design sketches (Section 5.3.3). This table was skipped in the original version due to space constraints.
1 Introduction
TABLE I: We found that simply presenting the table below (i.e., three heuristics along with explanations and examples) to participants at the beginning of the study can enhance the coverage of privacy-related design decisions in sketches, reduce the mental effort required for creating sketches, and improve the readability of the final sketches.
Heuristic Explanation Examples
Device-Based Data Flow Indicate the devices involved at each stage of the data flow to show how data moves from one point to another Capturing photos on a camera; saving user profiles on the server…
Stakeholder Interactions with Data Flow Show each individual or group’s interactions with the data flow, including involvement and privacy-related choices Manager authorizes AI modeling with user data; data scientist analyzes user’s activities…
Multi-Layered Representation Provide an overview of the privacy design, and then separate this from more detailed privacy considerations Displaying “storing data” in the overview, with details like “for 5 years with encryption” positioned in a separate layer (e.g., a different area within the sketch)
Figure 1: Example sketches for the Afterlife Chatbot scenario (see Table II), with sketches from participants exposed to heuristics (Table I) on the right and those without on the left. Heuristics-guided sketches (right) are more structured and comprehensive. The raw sketches (top) have been digitized (bottom) for improved readability.
Designing privacy in data practices is a collaborative process where data practitioners need to frequently communicate key privacy design concepts to others, such as how data is collected, used, shared, and how these data flows interact with users and other stakeholders [1, 2]. A few recent studies reveal that experienced data practitioners often use sketches, either on paper or with digital tools, to facilitate communication around privacy design concepts [3, 4].
Currently, there is no standardized approach to privacy sketching; instead, practitioners tend to improvise in an ad-hoc manner, combining multiple types of diagrams—such as data flow diagrams, use case diagrams, and component diagrams [4]. This ad-hoc privacy sketching resembles user experience (UX) sketching decades ago, where only a few UX designers could sketch the UX design concepts effectively based on unstructured intuitions and artistic skills [5]. Since then, HCI researchers have developed multiple structured frameworks to lower the barriers for UX designers to sketch user experiences [6].
In this paper, we draw inspiration from UX sketching and hypothesize that structured frameworks can help lower novice data science students’ barriers to creating high-quality privacy sketches for given data practice scenarios. To explore the potential framework, we conducted three studies with 54 unique participants across three U.S. universities, with no participant overlap between studies.
We began with a need-finding study to identify the challenges novice students face in communicating privacy designs (Section 3). We recruited 12 participants, asking them to consider two data practice scenarios, sketch their privacy design solutions on paper, and explain their designs using their sketches. This study revealed three key challenges: (1) novice students lack the appropriate vocabulary to sketch privacy designs, (2) participants struggle with planning and organizing the sketching space, and (3) the constraints of the sketching process can limit their ability to think expansively about design solutions.
We then adopted a human-centered design approach [7, 8] to guide the method’s development (Section 4). We iteratively prototyped a teaching method, tested it with a new group of users, gathered feedback, and refined the method iteratively. We explored three methods, ranging from a heavy-lifting approach with detailed digital diagramming to a lightweight method offering only three heuristics. Interestingly, the most lightweight method, which simply presents participants with a table of three heuristics (Table I), proved highly effective in improving the quality of novice students’ sketches despite its simplicity.
We conducted a detailed experiment to validate the effectiveness of the heuristic-based approach (Section 5). We recruited a total of 24 participants, with each participant creating a few sketches and interpreting the sketches from other participants. We found that participants who received heuristic-based instructions were able to cover 30% more of the privacy-related decisions in their sketches than the participants without exposure to these instructions. The final sketches from the participants who received heuristic-based instruction are also more readable, with an increase of 77% in interpretation accuracy.
Our primary contribution is a heuristic-based approach that lowers the barriers for novice students to create high-quality privacy sketches, along with insights gained from the iterative development process. Our exploration will help researchers understand how to teach data science students to communicate privacy designs. To the best of our knowledge, this is the first study to systematically explore methods for teaching novice students how to sketch privacy designs.
2 Related Work
This project builds on ideas from four key areas: (1) user experience and software design sketches, (2) education, (3) diagrams, and (4) developer support for privacy design.
2.1 User Experience and Software Design Sketches
Sketching has proven to be an efficient approach for exploring, communicating, and iterating on UX design ideas [9] and software design concepts [10, 11]. Sketching allows designers to rapidly visualize and refine user interactions while intentionally ignoring detailed specifications in early stages [12, 5]. This abstraction facilitates a smooth progression from initial concepts to functional design [13]. These findings have informed the development of structured sketching practices, where designers use specific frameworks and conventions to ensure clarity and consistency in visual communication [14, 15].
We hypothesize that an analogous approach can be applied to privacy design. Similar to how UX designers use sketches to map user journeys and identify pain points [9, 5], privacy practitioners can use sketching to illustrate how data is collected, processed, and shared and how the data flows interact with stakeholders.
2.2 Education for Privacy Design
Current privacy design education primarily focuses on privacy concepts [16] and design methodologies [17, 18], with most educators using case studies [17, 19] as the main form of instruction. For instance, some classes include a privacy review procedure [20, 19], where students assess smartphone privacy notices, identify issues, and propose design improvements in brief essays. Organizations such as the National Institute of Standards and Technology (NIST) [21] and the International Association of Privacy Professionals (IAPP) [22] also provide privacy design education by offering standards, certifications, and online learning resources, which provide practical guidance and structure for both academic and professional learning. Additionally, educators have expanded the instruction of privacy design to a broader audience, including industry professionals and policymakers [23]. In contrast, our work aims to complement existing curricula by exploring methods to teach data science students how to sketch privacy designs, equipping them with privacy design communication skills.
TABLE II: Task scenarios used in our study. We initially crafted 14 privacy-related scenarios from media reports [24, 25, 26, 27, 28, 29, 30, 31], research papers [32, 33, 3, 34], and product introductions [35, 36, 37, 38, 39, 40, 41, 42], then selected four scenarios based on participant familiarity, scenario complexity, and domain diversity.
# Scenario Description
1 Online Meeting Attention Tracking [25, 3] Designing an Attention Tracking feature for an online meeting app. The feature captures attendees’ focus and generates attention scores, enabling hosts to monitor engagement levels during meetings.
2 Financial Risk Management [39, 40, 41] Designing a feature for a bank’s mobile app to enhance risk management, including capabilities like anomaly detection and client credit assessments.
3 Afterlife Chatbot [24, 36, 37] Designing an “Afterlife Chatbot” service that allows clients to record video biographies. After the client passes away, their family can interact with the chatbot to preserve their memory.
4 Sensitive Image Detection [42, 38, 26] Designing an automated content moderation system for an image management service to detect inappropriate or harmful content in users’ uploaded pictures.
2.3 Diagrams for Privacy Design
Practitioners often use a combination of diagrams to communicate privacy considerations in system designs [3, 4, 43]. Research has shown that these diagrams help developers visualize system architecture by simplifying complex design elements, making it easier to communicate and design privacy-aware applications [4]. For example, when tasked with communicating the privacy design of an IoT application for diabetes treatment and monitoring, developers used Data flow diagrams (DFDs) [44] and Unified Modeling Language (UML) [45] to break down the flow of sensitive health data into manageable components, with annotations clarifying the encryption mechanism [4].
While existing studies shed light on the use of diagrams for communicating privacy considerations in system design, their focus has been on professional developers. However, how to teach these diagramming techniques to novice data science students remains unclear. Our work aims to address this gap by using sketches as the lens to investigate the specific challenges students face when communicating privacy design through diagrams and provide insights on targeted guidance to address these difficulties.
2.4 Developer Support for Privacy Design
Prior research has explored many approaches to support experienced developers in privacy design [46, 47]. For instance, some studies have offered insights into how to help developers make privacy design choices by identifying the common challenges in understanding and implementing privacy considerations [48, 49]. Additionally, various tools have been developed to support developers by embedding privacy practices directly into development workflows. For example, Coconut, an Android Studio plugin, helps developers create privacy-friendly apps by requiring privacy annotations and assisting them in organizing privacy-related details [50]. Similarly, PARROT helps developers create privacy-aware IoT applications by providing interactive privacy annotations and guidance [4].
However, existing studies focus on supporting experienced developers with existing knowledge of privacy. Little attention has been given to helping novice data science students learn to communicate privacy design. Our work bridges this gap by providing targeted support to help novices develop the skills needed to effectively communicate privacy design.
(a) Afterlife Chatbot (by N3).
(b) Financial Risk Management (by N6).
(c) Online Meeting Attention Tracking (by N10).
(d) Financial Risk Management (by N11).
[…]
3 Why is Sketching Privacy Challenging?
We began with a need-finding study [51] to identify barriers for data science students in sketching privacy designs.
Participants. To avoid priming, we advertised the study as a “data science experiment design study” rather than one specifically about privacy through social media and mailing listings across three U.S. universities. We recruited twelve students (seven identified as female, five as male) aged 19 to 24 (Mean = 21.7 years, SD = 2.1 years). Each participant was compensated for their time with a \$10 gift card. The sample included five graduates and seven undergraduates. We used the ACM Data Science Task Force’s definition of Data Science [52] to determine participants’ eligibility.
We asked participants, “Do you have any knowledge of privacy-related concepts? Please list” in a pre-screening survey to collect information about their privacy knowledge background. Among the 12 participants, 6 answered “No”; 3 listed relevant privacy design terms (e.g., privacy by design) they learned through research or course but lacked hands-on experience with privacy design; 3 provided responses unrelated to privacy.
Method. In each study, we randomly presented students with brief descriptions of two data practice scenarios (Table II). For each scenario, we asked them to consider the data practice, sketch relevant privacy designs on paper, and explain their designs to the researchers using their sketches. To understand the fundamental barriers they face, we did not impose time limits, allowing participants to sketch freely until satisfied. After the sketching session, three authors reviewed the participants’ sketches and asked additional questions to clarify concepts in their sketches. The authors also inquired whether they utilized any strategies or encountered any difficulties throughout their sketching process.
On average, participants took approximately 25 minutes (SD = 5.1 minutes) to complete a sketch. We summarized key insights after each study and identified data saturation [53] by the eighth study, with no new insights emerging. We then concluded with four more participants [54].
Findings. We made three main observations. Concepts in privacy design are abstract, and most students lack the appropriate vocabulary to sketch these concepts effectively. The process of sketching privacy is the process of discovering reusable visual vocabularies. Interestingly, most final sketches shared common components (Figure 2), such as stakeholder and device, but participants were unaware of these initially. Participants often started with an arbitrary line drawing and slowly recognized the basic vocabularies through iterations, and often re-drawn the sketch multiple times near the end using a few visual components discovered in the process. For example, when designing the Afterlife Chatbot service (the sketch without guidance in Figure 1), N10 initially focused on “client,” “record,” “video” and “family” to represent an outline. However, she quickly felt lost, unsure who would manage and process these videos to build the chatbot. She also tried adding the term “server” to indicate where the data would be stored and used. Besides, recognizing that not everyone could view the original videos and that only family members could use the service, she included “authorization” to suggest access control. In the end, her sketch became a mix of arrows, lengthy text, and icons as placeholders for unspecified roles.
Sketching privacy requires many iterations, and participants have difficulty planning the space in advance. A simple data practice can involve many low-level privacy-related design decisions, and many of them are interdependent. Participants often complained that the compact positioning (e.g., Figure 2) makes it hard to “read through the sketch” (N11). To make the problem worse, most participants need to iterate multiple rounds to have a decent design since it is hard to have a clear big picture of the privacy design before starting to sketch and think about the data practice. Participants reported that it is hard to “make targeted edits” (N2) and complained that sketching on paper lacked the flexibility to drag-and-drop for content rearrangement. In the end, participants either re-started a sketch to “re-plan the layout” (N6), squeezed updates into “any available blank space” (N2), or even “abandoned the idea to iterate” (N4). These challenges highlighted the need for better guidance and prompted us to explore teaching methods using digital tools (e.g., PC or iPad) to edit their already sketched content.
The limits of the sketches mean the limits of privacy designs. We encouraged participants to sketch while exploring the design space, but an unintended consequence emerged: they often stopped considering new design possibilities once they ran out of paper space. As a result, many sketches included surprisingly detailed components but overlooked the broader context of data practices. In the wrap-up interview, participants realized they “should include them but forgot” (N1). Others found it physically demanding to “resketch the components” (N2) already created in the previous steps. They expressed a desire for a “template” (N8) to guide them on how to add details to their designs and allow for “copy-paste” (N2) to reuse.
TABLE III: We iteratively experimented with three methods to teach data science students sketching privacy designs. This table summarizes the design changes and experimental observations of each method.
Teaching Method Participants Design Changes Interface Observations
Object-Oriented Diagramming T1-T6 (1) Reusable components (DataAction and Stakeholder classes); (2) Predefined attributes with preset options Web-based application (1) Learning curve associated with the tool structure; (2) Constrained design options due to incomplete attributes; (3) Appreciation for the re-usable components; (4) Preference for hand-sketch tools with copy-paste support
Vocabulary-Based Sketching T7-T11 (1) Visual vocabulary (e.g., “ellipse” for stakeholder); (2) Free text descriptions Tablet-based sketchboard (1) Appreciation for tablet-supported features; (2) Excessive attention to visual symbols; (3) Preference for flexible design guidelines over strict vocabulary; (4) Distraction from the overall design due to focus on details
Heuristic-Based Sketching T12-T18 Three heuristics: (1) Device-based data flow; (2) Stakeholder interaction with data flow; (3) Multi-layered representation Tablet-based sketchboard Preference for sketching with heuristics rather than vocabulary
4 Teaching Methods Experimentation
Design Goal. Our goal is to develop a method that enables data science students to create high-quality privacy sketches with minimal training. A high-quality sketch should meet two criteria: (1) it includes sufficient detail while covering the key components of the privacy design, and (2) it is clear and easy to interpret, allowing the audience to quickly understand the privacy concepts presented.
Experimentation Method. We employed an interactive prototyping method [7, 8] to guide the development of the teaching method. Interactive prototyping is a design method where users engage with evolving, often low-fidelity prototypes—such as sketches, paper interfaces, or mockups—in realistic scenarios while designers simulate system behavior, observe user interactions, offer feedback, and iteratively refine the design before the actual system is built. We iteratively prototyped a teaching method, tested it with a new group of users, gathered feedback, and refined the method. Since previous research found that 80% of usability problems are detected with four or five subjects and the most severe problems are likely to have been detected in the first few subjects [55], we experimented with each method using a relatively small participant sample (5-7 for each iteration).
Table III summarizes the process and the insights we obtained through the process. In this section, we will describe the three methods we explored, our observations, and the trade-offs associated with each. We recruited all participants in this experiment through the same approach as the need-finding study in Section 3. To maintain the integrity and independence of our findings, we also ensured there was no participant overlap across any of our studies, including future ones. Each participant was compensated for their time with a \$10 gift card.
(a) Stakeholder class.
(b) DataAction class.
[…]
4.1 Object-Oriented Diagramming
Our first attempt, object-oriented diagramming (OOD), is inspired by the object-oriented programming (OOP) principle in software engineering [56, 57].
4.1.1 Design
OOD has two key design ideas. First, we organized the privacy sketch based on the concept of objects, which can contain properties and methods (Figure 3) and support inheritance, similar to objects in OOP. We defined two base classes, DataAction and Stakeholder, along with their attributes. DataAction has four sub-classes (Collect, Store, Process, and Access). These subclasses share two common attributes, data and device, which indicate the direction of data flow (e.g., a photo moving from a camera to a server). Each subclass also has its specific attributes. Figure 3 presents the other class Stakeholder, which includes a name attribute for free text input. It consists of two methods: Involve and Decide. Involve denotes a stakeholder performing a DataAction (e.g., an admin storing data), while Decide represents making a choice (a binary parameter of the method, such as giving consent on data collection or revoking access).
Second, we offered preset attribute options to help novice students explore the design space systematically (Figure 3 and 4). Instead of describing each DataAction in natural language, students could select predefined attribute options. For example, when specifying data source to Collect, students could choose ‘direct’ for newly generated data or ‘secondary’ for pre-existing data.
4.1.2 Tool Support
We implemented our solution using Figma [58], a web application for UI design. This diagrammatic design employs three primary visual elements (Figure 3): rectangle, ellipse, and edge. Rectangles prompt students to choose from four predefined subclasses (Collect, Store, Process, and Access) with preset attribute options. Ellipses allow students to specify the stakeholder’s role with custom text. Edges depict relationships: a horizontal arrow between two rectangles indicates a sequential process, while a vertical connection from an ellipse to a rectangle represents either the stakeholder’s involvement (line) or a choice (arrow) that impacts the data flow.
4.1.3 Experiment
We first tested this approach with two undergraduate and two graduate data science students. We aimed to compare the object-oriented diagrams’ usability and users’ cognitive load with the hand-sketch method used in the pilot study (Section 3). Each participant created a privacy design using each approach, and we counterbalanced the order of approaches and randomly assigned two scenarios to each participant (Table II).
Because the prototype is low-fidelity—for example, it does not comprehensively enumerate all properties and methods for each object—we encouraged participants to interact with the experimenters to ask questions and clarify any uncertainties during the session. We observed that all four students experienced significant cognitive burdens when using OOD, which may have been partly due to the limited fidelity of our prototype. To address this, we recruited two additional privacy experts through personal connections to compensate for the limited fidelity—a common practice in interactive prototyping [7].
TABLE IV: In Vocabulary-Based Sketching, we incorporate visual components and use free text descriptions to replace the previous attribute structure. For testing, we presented participants with this worksheet, including each component’s name, symbol, and examples. Additionally, we specify the source and target nodes applicable to each edge component.
Vocabulary Symbol (Source, Target) Examples
Stakeholder User; admin; data scientist…
[…]
Involvement in the Data Flow (Stakeholder, Data Action) See examples above (i.e., performing the action indicates the stakeholder is involved in the data flow)
Choice on the Data Flow (Stakeholder, Data Action) User provides consent for data collection; manager authorizes AI modeling with user data…
Step Procedure between Data Actions (Data Action, Data Action) After collecting the data, proceed to save it…
4.1.4 Observations
All participants completed their hand sketches (Mean = 25.4 minutes, SD = 4.3 minutes), and our observations aligned with from the pilot study (Section 3). However, participants experienced a significant cognitive burden when using the diagramming approach. Only one privacy expert completed the privacy design in 23.1 minutes, whereas the remaining five participants required more time, ranging from 35 minutes to over an hour. Most participants felt their privacy designs were incomplete and eventually reached a point where they couldn’t make further progress. Four participants chose to terminate the task without satisfaction after it extended beyond an hour, explaining that they “became lost” (T3) and were “unsure how to improve the design” (T5), for several reasons:
Participants expressed the need for more time to become familiar with the web interface and the Object-Oriented concepts. For example, one user mentioned spending a significant amount of time “navigating each component and its attributes” (T1) to understand how it fits into their design, which they “wouldn’t need to do with a sketch” (T4).
Next, participants raised concerns about the constrained design options provided to them. For instance, all participants found the preset options of attributes limiting, including T2, who completed her diagram (Figure 4). “I wanted to indicate the input as ‘original photo’ and output as ‘labeled photo with detection results’ in the processing step. However, the only option for me was ‘photo,’ which is hard to infer if someone was reading my design.” Others also expressed that they prefer to “use plain text to describe” (T1) their designs. T5 complained, “This rigid structure made me feel like I was designing within a box.”
On the other hand, participants appreciated the clarity of visual shapes and the efficiency offered by reusable components. Many noted that preset elements significantly streamline the sketching process, as they found such elements “impractical for on-paper sketching” (T3). Participants also expressed a strong desire to use tablets for editing efficiency. “It would be more natural to work on an iPad - I could easily duplicate and reuse components” (T2).
4.2 Vocabulary-Based Sketching
Building on the previous observations, we refined our teaching method in design and tool support.
4.2.1 Design
A Vocabulary of Privacy Design Components. We relaxed the hierarchy requirement of the object-oriented design and simplified the vocabulary set (Table IV). For example, we eliminated the subclasses of DataAction and defined a Data Action node represented by a rectangle. Specifically, as one of the common components in the pilot study (Section 3), <DEVICE> was designed as a free text handwriting input to specify where the data action occurs. Additionally, the two methods (Figure 3) of the Stakeholder class and the process between data actions were represented as three types of edges connecting nodes.
Free Text Descriptions. To address the concern of constrained design space, we removed the design of predefined attributes. Instead, we allowed participants to use free text to describe privacy-related details.
4.2.2 Tool Support
As our last experiment suggested, rather than a web-based interface, users preferred hand-sketch tools with copy-paste functionality, where they could reuse the created components. We adapted our teaching method to a tablet-based sketch board. For example, students could use an iPad with a stylus to create their privacy designs on a sketch board app such as Notability [59].
4.2.3 Experiment
We tested this approach with three undergraduate and two graduate data science students. Due to frequent difficulties in completing the diagrams (Section 4.1.4), we did not ask them to test the Object-Oriented Diagramming in this round. We also asked each participant to sketch only one randomly selected scenario (Table II). To support participants’ learning process, we provided a worksheet (Table IV) with vocabulary, symbols, and usage examples. We guided them through its content before they started the design tasks and encouraged them to add free-text privacy details. For example, they may specify whether encryption was applied when saving user data.
4.2.4 Observations
All participants finished their privacy design sketches (Mean = 18.6 minutes, SD = 2.7 minutes) and took advantage of the sketch board’s editing features to erase, copy-paste, drag, or resize their created content (Figure 5). However, they raised several concerns about this vocabulary-guided approach.
First, participants reported focusing too much on the visual components, which they felt was “not very intuitive” (T7). They expressed needing to “constantly stop and think about how to use each element” (T10). Four participants wanted flexible guidance, with T8 saying, “Just give me a simple guideline, like including the stakeholder’s engagement, so I can focus on the design without worrying about using rectangles or ellipses.” (T8).
Second, participants struggled to balance detailed privacy considerations while maintaining a design overview (Figure 5). For instance, T10 noted, “When I was deciding how long data should be retained on the server, I lost track of the next data action I wanted to sketch.” Three participants desired an “extra page” to capture such details, with T9 suggesting, “An extra page for a Data Action node could include data retention periods, so I could simply edit that page if I wanted to update it from 1 month to 1 year.”
[…]
4.3 Heuristic-Based Sketching
Based on the previous findings and teaching method designs (Section 3 and Table III), we moved toward a more lightweight solution that explores heuristics as the main approach to teach students to sketch privacy designs.
4.3.1 Design & Tool Support
We further simplified our method by removing constraints on vocabulary and visual representations, allowing participants to focus on the content of the design. This modification led to three heuristics (Table I) for sketching privacy designs.
[…]
4.3.2 Experiment
We tested the Heuristic-Based Sketching against Vocabulary-Based Sketching with four undergraduate and three graduate data science students. Each participant sketched privacy designs using both approaches, and we counterbalanced the order of approaches and two scenarios presented to them (Afterlife Chatbot and Financial Risk Management, see Table II). We chose these two scenarios because previous participants found these topics interesting, which enhances participants’ engagement.
To support their learning process, we included a warm-up session before each task based on feedback from previous participants (Section 4.2.4). We introduced an example scenario (Online Meeting Attention Tracking, see Table II) and guided them through the corresponding worksheet (Table I or IV). Participants then sketched the example scenario and asked any questions only during warm-up.
4.3.3 Observations
All participants completed the warm-up within 20 minutes (Mean = 18.5 minutes, SD = 1.2 minutes) and reached satisfaction with each design within 15 minutes (Mean = 13.6 minutes, SD = 2.9 minutes). First, we observed a significant reduction in completion time with the vocabulary-based approach (14.3 v.s. 18.6 minutes in Section 4.2.4), as participants reported that the warm-up session helped them become familiar with the method. Additionally, there was a clear preference for heuristics over vocabulary-based design. For example, participants who started with heuristic-guided sketching often intentionally applied the multi-layered heuristic when sketching vocabulary-based designs. However, we did not observe a tendency to re-apply visual vocabulary to heuristic-based tasks. T13 noted, “I felt freer to move beyond ellipses and boxes and to organize my ideas in a more trackable way.”
5 Evaluation
Figure 6: Each evaluation study included a warm-up session, two creation tasks (followed by an annotation process of privacy-related design decisions), two interpretation tasks, and a post-task interview. During annotation, participants also orally explained the content of each design decision. They completed a NASA TLX survey after each task.
This section presents a detailed experimental evaluation of the heuristic-based approach. Our results indicate that this approach helps data science students create high-quality privacy sketches quickly with reduced mental effort. Participants also reported that the guided privacy sketches were more readable than the baseline sketches.
5.1 Study Overview
Participants. We recruited 24 participants in the evaluation through the same approach as previous studies (Sections 3, 4). Our final group included 10 undergraduates and 14 graduate students. In the pre-screening privacy knowledge survey (same as Section 3), 14 participants answered “No” to the privacy knowledge question; six gave responses unrelated to privacy, and only four provided relevant privacy design terms but lacked experience in privacy design.
Among our participants, 10 (41.7%) identified as female, and others identified as male. Twenty (83.3%) participants were aged between 18 and 24, while others were between 25 and 30 (Mean = 22.0 years, SD = 2.3 years). Each participant received a US \$15 Amazon gift card as compensation.
Study Procedure and Apparatus. We followed a between-subject study design. Participants received either vocabulary-based or heuristic-based guidance for their tasks. They completed tasks on an iPad with a stylus using a sketch board application (e.g., Notability [59]).
Figure 6 illustrates the study procedure. Each study began with a warm-up session (20-minute limit) that included an explanation of the worksheet (Table I or IV) and sketching privacy design for a sample scenario Online Meeting Attention Tracking (Table II). Researchers observed the process and answered questions as needed.
Then, the study continued with two creation tasks in which participants sketched designs for two scenarios (i.e., Afterlife Chatbot and Financial Risk Management, as shown in Table II) and annotated their privacy design decisions in the sketches. The participants were randomly assigned to a scenario (Table 5.2.1), and upon completion, they would complete another. Each task has a 15-minute time limit. These time limits were decided based on the task completion time in the last experiment (Section 4.3.3).
Then, each participant was asked to interpret two sketches created by previous participants (Table 5.2.4). Finally, we interviewed participants about what hindered or assisted them in their creation and interpretation tasks.
We took notes during each study session and summarized the key insights after each session. We observed saturation (i.e., no new insights emerged) [53] after the twentieth study. We then stopped participant recruitment and concluded the evaluation study with four more participants[54]. On average, the study session lasted 90 minutes.
During analysis, we performed Mann-Whitney U tests [63] to compare different scenario orders within each condition and found no evidence of ordering bias across all measures reported in Sections 5.2.3 and 5.3.3. For example, within the vocabulary group, we assessed the first creation task’s mental demand scores between participants who began with scenario F and those who began with scenario A.
In the following sections, we will denote the Vocabulary-Based Sketching as “Vocabulary” and the Heuristic-Based Sketching as “Heuristic.”
Annotating Visual Sketches. Comparing high-dimensional sketches is challenging. To address this, we developed a codebook (Appendix Table XI) to annotate privacy-related design decisions covered in the sketches. We then used the annotated decisions to assess the design coverage in sketch creation and the communication effectiveness between the sketch creators and interpreters (Sections 5.2, 5.3).
Since participants often cannot broadly explore the design space (Section 3), we cannot derive the potential design decisions entirely based on empirical observations. Instead, we began by collecting privacy-related design decisions from prior literature [64, 32, 65, 66, 67, 68] to systematically capture the unique design decisions reflected in the sketches. Two researchers then collaboratively created an initial codebook for analysis. They then independently applied the codebook to annotate the sketches from previous studies (Sections 3, 4) and discussed potential modifications, deletions, and extensions to the coding scheme. After refining the scheme, all these previous sketches were re-coded using the updated framework. Given the subjective nature of interpreting sketches, we considered only codes independently validated by both researchers to ensure reliability.
5.2 Creation Tasks
5.2.1 Data Collection
We counterbalanced the presentation order of scenarios (Table 5.2.1). In each task, we asked participants to read the scenario description (Table II) and sketch a privacy design. They could inform us if they felt satisfied with their sketch, i.e., having included as many privacy-related details as possible, before the time was up. We recorded their completion time and asked them to fill out a NASA TLX survey [69] after each creation.
We presented participants with a list of privacy-related design decisions (Table XI) and asked them to describe each decision they had included, annotating the corresponding content on their sketches. We didn’t set time limits for this process. Participants could skip any decisions they felt were not covered but could not modify their original designs.
TABLE V: Creation task schedule. We counterbalanced the presentation order of the task scenarios (F: Financial Risk Management; A: Afterlife Chatbot) for both conditions (Vocabulary-guided and Heuristic-guided sketching).
Condition Scenario Creator
1st 2nd
Heuristic F A P1 P3 P5 P13 P15 P17
A F P7 P9 P11 P19 P21 P23
Vocabulary F A P2 P4 P6 P14 P16 P18
A F P8 P10 P12 P20 P22 P24
5.2.2 Data Analysis
There is no perfect sketch for the creation task, as participants may have different designs for the same scenario, which may evolve during sketching. So, we use the codebook (Table XI) to systematically capture the unique design decisions reflected in each sketch. To assess the quality of the sketches, we documented the privacy design decisions annotated by participants, counted the total number of decisions included in each sketch, and calculated the coverage percentage of each design decision across all sketches in each condition.
Next, we conducted a thematic analysis [70] of post-task interview transcripts and notes we took during study. First, two researchers independently coded eight transcripts (four from heuristic group and four from vocabulary group). After discussing and incorporating codes, we created a codebook that the two researchers agreed on. Using this codebook, the researchers divided the remaining transcripts, each coding eight transcripts per condition. Like many other qualitative analyses in S&P research [71, 72], the two researchers discussed and resolved coding conflicts in several weekly meetings. In this analysis, as we prioritized identifying emerging themes, we did not calculate the inter-rater reliability (IRR) to seek theoretical agreement [73].
TABLE VI: In creation tasks, heuristic-guided participants spent less time sketching until satisfaction and included more privacy-related design decisions in their sketches. Format: mean $\pm$ standard deviation. We also highlight the higher value between two conditions.
Task Order Vocabulary Heuristic
Time (min) # Decisions Time (min) # Decisions
First 14.89 $\pm$ 2.86 8.4 $\pm$ 3.0 12.70 $\pm$ 0.86 10.7 $\pm$ 2.8
Second 14.70 $\pm$ 3.71 9.3 $\pm$ 3.1 10.36 $\pm$ 0.98 12.3 $\pm$ 1.8
All 14.80 $\pm$ 3.23 8.9 $\pm$ 3.0 11.53 $\pm$ 0.94 11.5 $\pm$ 2.4
5.2.3 Quantitative Results
We found that heuristics could help students sketch privacy designs more effectively. Their designs also demonstrated higher quality.
Improved Sketching Efficiency. As shown in Table VI, heuristic-guided participants reached their satisfied designs faster (average of 11.53 v.s. 14.80 minutes) and covered more privacy-related design decisions (average of 11.5 v.s. 8.9 per sketch), compared to the vocabulary group.
Besides, as they progressed to the second creation task, the heuristic-guided participants reported more positive responses to the NASA TLX questions (Table VII) than the first task. These improvements included a significantly reduced mental workload shown by the Wilcoxon Signed-Rank test [74] ($p~{}\textless~{}0.05{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},r=0.70}$), as well as decreases in physical, temporal, and effort demands. The heuristic group also noted increased self-perceived performance and reduced frustration, a trend not observed in the vocabulary group.
Broader Coverage of Privacy-Related Design Decisions. We then compared the coverage percentage of design decisions between conditions (Table VIII). On average, each decision was included in 76.7% of the heuristic-guided sketches, compared to 59.2% of the vocabulary group. Specifically, heuristic group could more frequently cover decisions such as Stored Data, Processing Input, Processing Approach, Accessed Output, Access Approach, and Choice Impacts. This difference was statistically significant ($p~{}\textless~{}0.01{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},r=0.69}$) under the Wilcoxon Signed-Rank test [74].
TABLE VII: In the second creation task, heuristic-guided participants reported more positive responses than they did in the first task, a trend not observed in the vocabulary group. We present NASA TLX results (scale of 1 to 5) as “median (mean $\pm$ standard deviation),” and highlight second task responses if they are more positive than the first. “” indicates statistical significance ($p~{}\textless~{}0.05$) under Wilcoxon Signed-Rank test. $\downarrow$” denotes that a lower value is a more positive outcome.
Condition Task Order Mental Demand $\downarrow$ Physical Demand $\downarrow$ Temporal Demand $\downarrow$ Performance $\uparrow$ Effort $\downarrow$ Frustration $\downarrow$
Vocabulary First 3.5 (3.17 $\pm$ 1.03) 4.0 (3.65 $\pm$ 1.23) 4.0 (3.33 $\pm$ 0.98) 3.0 (3.00 $\pm$ 0.74) 4.0 (3.83 $\pm$ 0.39) 2.0 (2.50 $\pm$ 0.90)
Second 3.5 (3.33 $\pm$ 0.78) 4.0 (3.81 $\pm$ 1.31) 3.5 (3.17 $\pm$ 0.94) 3.0 (2.92 $\pm$ 1.08) 3.5 (3.50 $\pm$ 0.80) 3.0 (2.67 $\pm$ 0.89)
Heuristic First 4.0 (3.67 $\pm$ 0.89) 3.5 (3.48 $\pm$ 0.97) 4.0 (3.58 $\pm$ 0.79) 3.0 (2.92 $\pm$ 0.67) 3.5 (3.58 $\pm$ 0.67) 3.0 (3.00 $\pm$ 1.04)
Second 2.5 (2.92 $\pm$ 1.24) 3.0 (2.97 $\pm$ 0.82) 3.0 (3.00 $\pm$ 1.41) 3.5 (3.25 $\pm$ 0.87) 3.0 (3.17 $\pm$ 1.19) 2.0 (2.33 $\pm$ 1.07)
5.2.4 Qualitative Findings
Based on the analysis of participants’ creation task behaviors and their feedback in post-task interviews, we identified following main findings.
Reduced Learning Curve. The guidelines provide a starting point and ease the challenge of beginning a sketch from scratch. For example, P15 shared that “The heuristic helps me to know what the design should roughly look like when I started to construct my ideas.” P7 added that “Including stakeholders prompts me to consider who is involved and who is taking action in data cases, which I believe is important when forming my ideas for sketches.”
Flexibility in Planning. Participants appreciate the Multi-layered Representation heuristic for helping them in planning their sketches in a structured manner. P7 explained, “The overview layer allows me to lay down the general ideas so I can worry about the details later.” The device-based data flow and multi-layered approach also gave them a “big-picture perspective” (P3), which helped P5 to “organize the system’s overall logic and focus on how data flows between different modules without getting distracted by other details.” Participants also highlighted that the separate layers enable them to “quickly update on details” (P9) and “easily to replicate existing content” (P11). Furthermore, as they were familiar with the heuristics after the first design, participants in the experimental group were able to quickly grasp and re-apply the workflow to the second task.
Ease of Cognitive Load. After applying the heuristics, participants recognized that sketching was more manageable than anticipated, reducing their perceived effort and workload. P7 expressed that using the device annotation made “each part more distinctive and save effort if I want to adding new ideas to my sketch.” P19 initially worried about “how to show all the privacy details” but later found that “the layered approach helped me break down complex ideas into simple parts.” Despite the initial time pressure, P5 noted the benefit of sketching the overview layer first, explaining that it allows for “managing time more effectively by following the outline I planned out from the start.”
Enhanced Coverage of Privacy Details. The heuristics prompt participants to extend the scope of privacy details in their sketches, echoing the increased design decision coverage shown in Table VIII. First, the Stakeholder Interactions with Data Flow heuristic encouraged the participants to “consider human decisions and accountability with freedom” in a data practice (P11), which is “important for ensuring responsible use of data” (P7). P23 explained that the heuristic enabled thinking “beyond where the data travels in the system” and encouraged consideration of “what each party is doing with the data.”
Second, the Multi-Layered Representation heuristic encourages more detailed privacy designs. Participants noted that the separation of layers “makes space for me to fill in details” (P9) and that the flexibility to update individual layers enabled easy extension of their sketches, encouraging them “to iterate more times without friction” (P3).
As a result, compared to the vocabulary group, heuristic-guided sketches incorporated more interaction details beyond merely including “Choice” and “Involvement.” For example, as illustrated in Figure 1, these sketches featured interactive elements such as a chat window for access or a pop-up for making choices. This corresponded to the increased coverage (Table VIII) of the three decisions related to Choice & NoticeChoice Options, Choice Impacts, and Choice Notification.
Bias and Negative Consequence. Participants noted certain limitations with the heuristics during the sketching process. P1 remarked that the Multi-layered Representation “seems to be repetitive because I am filling similar information in both layers.” P7 observed that creating a sketch on the given case was “a systematic design” and expressed a desire for a more “complete structured framework to fill or follow.”
TABLE VIII: Heuristic-guided sketches could include privacy design decisions more frequently. Additionally, the design decisions within the heuristic-guided sketches were interpreted more accurately than those within the vocabulary group. For each design decision, we report its coverage frequency (“Coverage”) across all sketches in each group. For communication, we report the frequency of being interpreted (“Interpret,” including misinterpretations), along with precision, recall, and F1-score, all measured at the design decision level. For each metric, we highlight the higher value and indicate the statistical significance between the two conditions using the Wilcoxon Signed-Rank tests (${}^{*}p~{}\textless~{}0.05$, ${}^{**}p~{}\textless~{}0.01$, ${}^{***}p~{}\textless~{}0.001$), all showing large effect sizes ($r~{}\textgreater~{}0.5$).
Design Decision Vocabulary-Guided Sketches Heuristic-Guided Sketches
Coverage Interpret Precision Recall F1 Coverage Interpret Precision Recall F1
Collected Personal Data 100.0% 86.4% 83.3% 92.1% 87.5% 95.8% 87.0% 90.9% 100.0% 95.2%
Data Provider 87.5% 95.5% 70.5% 73.8% 72.1% 91.7% 95.7% 91.3% 95.5% 93.3%
Collection Purpose 87.5% 77.3% 27.8% 29.4% 28.6% 83.3% 73.9% 65.0% 76.5% 70.3%
Stored Data 62.5% 77.3% 34.4% 32.4% 33.3% 91.7% 78.3% 77.8% 77.8% 77.8%
Storage Approach 79.2% 63.6% 33.3% 42.9% 37.5% 91.7% 73.9% 78.9% 88.2% 83.3%
Post-Storage Action 58.3% 31.8% 10.0% 21.4% 13.6% 62.5% 60.9% 56.3% 64.3% 60.0%
Processed Input 54.2% 81.8% 29.4% 27.8% 28.6% 100.0% 78.3% 67.5% 75.0% 71.1%
Processing Output 79.2% 81.8% 55.3% 58.3% 56.8% 91.7% 65.2% 85.3% 96.7% 90.6%
Processing Approach 45.8% 50.0% 29.2% 31.8% 30.4% 83.3% 60.9% 66.7% 71.4% 69.0%
Accessed Raw Data 66.7% 54.5% 21.4% 25.0% 23.1% 50.0% 52.2% 59.1% 54.2% 56.5%
Accessed Output 41.7% 63.6% 58.3% 50.0% 53.8% 75.0% 60.9% 52.8% 67.9% 59.4%
Access Approach 50.0% 59.1% 31.8% 26.9% 29.2% 83.3% 69.6% 66.7% 62.5% 64.5%
Choice Options 37.5% 50.0% 45.5% 45.5% 45.5% 58.3% 52.2% 64.3% 75.0% 69.2%
Choice Impacts 16.7% 36.4% 14.3% 12.5% 13.3% 58.3% 30.4% 37.5% 64.3% 47.4%
Choice Notification 20.8% 18.2% 16.7% 25.0% 20.0% 33.3% 26.1% 75.0% 75.0% 75.0%
Mean 59.2% 61.8% 37.4% 39.7% 38.2% 76.7% 64.4% 69.0% 76.3% 72.2%
[…]
TABLE IX: Interpretation task schedule. For instance, P3 first interpreted a sketch of Financial Risk Management, which was created by P1 (heuristic group), and then interpreted a sketch of another scenario by P2 (vocabulary group). The “-” means no interpretation task due to the lack of a prior sketch, and three sketches (one by P23 and two by P24) were not interpreted because of no subsequent interpreters.
Scenario F A Interpreter Scenario A F Interpreter
Condition Heuris. Vocab. Condition Vocab. Heuris.
Creator - - P1 Creator P6 P5 P7
P1 P2 P3 P8 P7 P9
P3 P4 P5 P10 P9 P11
P11 P12 P13 P18 P17 P19
P13 P14 P15 P20 P19 P21
P15 P16 P17 P22 P21 P23
Scenario F A Interpreter Scenario A F Interpreter
Condition Vocab. Heuris. Condition Heuris. Vocab.
Creator - P1 P2 Creator P7 P6 P8
P2 P3 P4 P9 P8 P10
P4 P5 P6 P11 P10 P12
P12 P13 P14 P19 P18 P20
P14 P15 P16 P21 P20 P22
P16 P17 P18 P23 P22 P24
TABLE X: Participants perceived significantly lower workloads when interpreting heuristic-guided sketches than vocabulary-guided ones. We present their responses to NASA TLX survey as median (mean $\pm$ standard deviation), with results for heuristic-guided sketches highlighted. We annotate statistically significant improvements based on the Wilcoxon Signed-Rank tests (${}^{*}p~{}\textless~{}0.05$, ${}^{**}p~{}\textless~{}0.01$, ${}^{***}p~{}\textless~{}0.001$), all with medium to near large effect sizes ($r:0.3\sim 0.5$). $\downarrow$” indicates that a lower value is more positive.
Condition Mental Demand $\downarrow$ Physical Demand $\downarrow$ Temporal Demand $\downarrow$ Performance $\uparrow$ Effort $\downarrow$ Frustration $\downarrow$
Vocabulary 3.0 (3.41 $\pm$ 1.10)∗∗∗ 3.5 (3.66 $\pm$ 0.92)∗∗ 3.0 (2.82 $\pm$ 1.18) 3.0 (2.91 $\pm$ 1.15)∗∗ 3.5 (3.32 $\pm$ 1.04)∗∗∗ 3.0 (3.23 $\pm$ 1.07)∗∗∗
Heuristic 2.0 (2.29 $\pm$ 0.86)∗∗∗ 2.5 (2.61 $\pm$ 0.88)∗∗ 2.0 (2.17 $\pm$ 0.87) 4.0 (3.75 $\pm$ 0.79)∗∗ 2.0 (2.25 $\pm$ 0.85)∗∗∗ 2.0 (2.04 $\pm$ 0.81)∗∗∗
5.3 Interpretation Tasks
5.3.1 Data Collection
As shown in Figure 6, each participant completed two interpretation tasks, each with a 7-minute time limit. In each task, they received a randomized list of privacy-related design decisions (excluding the “Procedure” column in Table XI) and an unannotated copy of a sketch created by a prior participant. We then asked them to annotate and orally describe the privacy-related design decisions they identified. This annotation is similar to what they have done during the creation tasks (Section 5.2.1). Table 5.2.4 outlines the interpretation task order, where each participant interpreted two sketches of different scenarios: one created by a heuristic-guided participant and another from the vocabulary group. We maintained scenario order consistent with the creation tasks (Table 5.2.1) to avoid biases from the most recent scenario they had sketched. For example, after sketching for Afterlife Chatbot as his second creation, P3 would first interpret a sketch of another scenario. After each task, we recorded their interpretation time and asked them to complete a NASA TLX survey [69].
5.3.2 Data Analysis
To assess whether heuristic-guided sketches could better facilitate communication of privacy design, we measured the alignment between the creator’s and interpreter’s responses (including annotations and explanations) regarding each privacy-related design decision within a sketch. For each design decision, two researchers collaboratively compared both response versions and assigned a code to describe whether they aligned. They open-coded a subset of sketches (N=10, i.e., 15 codes per sketch $\times$ 10 sketches), conducted three rounds of discussions to reach a complete agreement, and generated an initial codebook. Then, two researchers coded the rest of the sketches independently. To ensure consistency, two coders discussed their codes regularly to reach an agreement and iteratively refined the codebook (e.g., an initial code “only one person gave a response” was decoupled into two codes “only the creator responded” and “only interpreter responded”). Finally, two researchers coded all the sketches using the final codebook (Appendix Table XII), with a Cohen’s kappa [75] of 90.1%, which reflects an “almost perfect agreement” [76].
We then quantified the effectiveness of communication on two levels. First, to measure whether heuristic-guided sketches lead to high-quality interpretations for each privacy-related design decision, we used the creators’ responses as groundtruth and calculated design decision-level precision (i.e., the proportion of interpretations that matched the groundtruth) as well as recall (i.e., the proportion of groundtruth responses that were accurately interpreted).
Second, to determine whether sketching with heuristics aids in identifying correct privacy design decisions across an entire sketch, we computed sketch-level precision (the proportion of interpretations that matched the creator’s responses within each sketch) and recall (the proportion of design decisions covered by the creator that was accurately interpreted in each sketch). Appendix Table XIII provides further clarification of these metrics.
The qualitative analysis shared the same procedure as described in Section 5.2.2, and we will report the findings of interpretation tasks in Section 5.3.4.
5.3.3 Quantitative Results
We found that heuristic-guided sketches are easier to interpret and could better facilitate communication between creators and interpreters.
Eased Workload of Interpretation. All participants finished their interpretation within the time limit. There was no significant difference in interpretation time (4.9 minutes for heuristic-guided sketches v.s. 5.1 minutes for vocabulary-guided sketches). However, participants consistently reported significantly lower mental, physical, temporal, and effort-related demands and reduced frustration when working with heuristic-guided sketches. They also perceived improved performance, as indicated by Wilcoxon Signed-Rank tests (Table X). Furthermore, participants identified more privacy design decisions in heuristic-guided sketches, with an average of 9.4 elements per sketch compared to 8.1 from vocabulary-guided ones. These results suggest that heuristic-guided sketches are more straightforward to interpret, decreasing workload while enhancing outcomes.
Improved Communication Efficiency. Table VIII presents the result of interpretation evaluation metrics at the privacy design decision level, which provides initial evidence for heuristic guidance’s capability to enable better creator-interpreter communication of privacy designs. First, heuristic-guided sketches achieved higher interpretation performance across metrics, with an increase of 31.6% in precision (i.e., increased fraction of correct interpretations) and 36.6% in recall (i.e., increased fraction of creator’s design decisions that were accurately interpreted).
At the level of a complete privacy design, heuristic-guided sketches demonstrated higher per-sketch interpretation performance, again with a higher precision (79.1% v.s. 44.6% for the vocabulary group), reflecting the proportion of matches among all interpretations in a sketch, and higher recall (71.1% v.s. 40.5%), representing the percentage of alignments among the creator’s all design decisions. These differences are all statistically significant according to the Wilcoxon Signed-Rank tests [74], highlighting the improved communication efficiency facilitated by heuristic guidance.
Figure 7: A heuristic-guided privacy design sketch for the Afterlife Chatbot scenario. It presents privacy details in parallel with the corresponding devices, stakeholders, and the steps taken, offering a clear logic for readers to follow.
5.3.4 Qualitative Findings
We made the following findings based on participants’ behaviors and feedback about their interpretation tasks.
Heuristics streamline the interpretation process. Participants from both conditions found heuristic-guided sketches to be more structured and easier to understand, largely due to the enhanced readability provided by the Multi-Layered Representation heuristic. P17, a heuristic-guided participant, appreciated how the multi-layered design allowed him to quickly grasp the overall idea of the privacy design at first glance, noting that “[reading] this sketch felt quite similar to [creating] my own sketch work.”
Similarly, P10, who did not receive heuristic instruction, also found the sketch (Figure 7) easier to follow, explaining that “the overview layer explains a clear logic of the design, such as where the privacy-sensitive procedure happens and who were involved in. However, the other [vocabulary-guided] sketch obviously falls short in this clarity.” She further expressed interest in “adopt[ing] this layered style in my future design sketches.” P4 also emphasized that separating the overview from details “enables a progressive manner of reading,” while P18 highlighted how “this multi-level sketch prevents my visual overload from digesting too much content in one place.”
Heuristics offer varying content density. Participants showed different preferences when retrieving information: some preferred more text for detailed understanding, while others favored less text for a smoother reading experience. The heuristic-guided sketches effectively accommodated both needs. P6 praised the multi-layered sketch for its support of effective information retrieval, saying, “first, I skim through the overview of the data flow to understand the big picture. If I want specific details, like the data involved in each step, I go to the next layer using the annotation labels.” P10 described the device box and overview layer as “pretty concise and simple,” while preferring more text for detailed understanding. P11 appreciated that the “detail layer has enough information for me to understand creator’s thought,” but felt that excessive text made interpretation time-consuming.
6 Discussion & Future Work
Connections to Established Practices. Professionals often draw on their existing skill sets when sketching for privacy. For instance, UX designers may use Customer Journey Maps, while software engineers rely on UML diagrams. We chose to develop a tailored method from scratch for two reasons. First, these skill adaptions only partially address key aspects of privacy design. For example, Customer Journey Maps focuses on stakeholder interactions. Second, most of our target audience—data science students—lacks formal training in both usability research and software engineering.
We have incorporated multiple relevant ideas in software engineering and UX research into our solutions, such as Data Flow Diagrams, Customer Journey Maps, and UX storyboards [77]. Future work may investigate whether our proposed heuristics can also enhance the practices of experienced professionals and how these heuristics complement their existing techniques.
Why do heuristics work in sketching privacy? Unlike rigid frameworks that limit creativity and restrict exploration [4, 44], the heuristic-based approach emphasizes flexibility and provides simple guidance, enabling students to deeply engage with privacy contexts while effectively organizing their ideas. Furthermore, the Multi-layer Representation heuristic helps students manage their sketches from a macro perspective, allowing them to approach designs strategically.
Sketches v.s. Diagrams. We experimented with both sketches and diagrams in this project. We found that complex diagram structures hindered the learning process (Section 4.1.4). For instance, when students used diagrams to illustrate interactions between stakeholders and data flows, they needed a thorough understanding of diagram vocabulary. This steep learning curve increased their cognitive load, making it harder to complete the privacy design.
In contrast, sketches mitigated these challenges by enabling more intuitive idea generation (Section 5.2) and communication (Section 5.3). This aligns with prior research highlighting sketching as a flexible and intuitive tool for ideation and communication [5, 78]. However, sketches lack the structured detail of diagrams, which is crucial for tasks requiring precision and depth. While sketches reduce cognitive demands, they may limit students’ ability to achieve the detailed understanding offered by structured diagrams. Future work could explore methods to help students effectively use diagrams to communicate privacy designs.
Sketching for Privacy Literacy. In this paper, our educational audience is data science students who are future data practitioners. One potential future direction is to push the technique to layperson [79], and explore if sketching privacy can help them form a more precise mental model [80] of the data practices.
High-fidelity Privacy Prototypes. Two students with UX/UI backgrounds noted that, although they were generally satisfied with their privacy designs, they felt uneasy about their inability to fully depict the visual aspects of the product’s interface. They also found that messy hand sketches contributed to difficulties in interpretation. Future work could explore higher-fidelity privacy prototypes, akin to high-fidelity UX prototypes.
Sketches and Threat Modeling. Our current study focuses on facilitating the communication of privacy design concepts through sketches. Future research may further explore the benefits of sketches in other security and privacy applications. For example, participants may develop a stronger awareness of privacy risks during sketching [81, 82, 83, 84]. The improved communication enabled by privacy sketches may enhance the collaborative decision-making process [1]. Conventional threat modeling approaches [85, 86] often rely on static tables or data flow diagrams (DFDs) to define technical scope and decompose applications. Compared to tables and DFDs, sketching may better capture dynamic aspects of a system, such as stakeholder interactions and contextual nuances that are difficult to convey textually.
7 Limitations
Ground Truth for the Creation Task. The creation task has no ground truth. While our codebook (Table XI) provided a structured approach for labeling and interpreting data, it may have constrained the granularity of our analysis. By relying on predefined categories, we may have overlooked subtle contextual differences or nuances in participant responses.
Participant Pool Limitations. Our participant pool consisted primarily of students from four-year U.S. universities, which introduced limitations in both sample diversity and size. In particular, some participants had prior exposure to privacy-related concepts, which may not reflect the perspectives of novice users who engage with privacy features with little to no foundational knowledge. Future research may expand recruitment to include participants from community colleges and technical bootcamps, capturing a broader range of educational backgrounds and user experiences.
8 Conclusion
This paper explores methods for teaching data science students to sketch privacy designs. Through a need-finding study (N=12), we identified three challenges students encounter when sketching privacy designs: (1) difficulty sketching a complete data flow, (2) omission of stakeholder interactions with the data flow, such as user consent, and (3) failure to make quick updates without disrupting the overall design. To address these challenges, we iteratively developed our teaching approach, ultimately leading to three heuristics: (1) Device Annotation, (2) Stakeholder Interaction with Data Flow, and (3) Multi-Layered Representation. Our between-subjects experiment (N=24) demonstrates the effectiveness of these heuristics in improving sketching quality and facilitating communication of privacy design. Future privacy education and privacy design toolkits could take our findings into consideration.
9 Research Ethics
This research received approval from the Institutional Review Board (IRB) of our institution. We conducted all studies using Zoom (approved by our institute) and used Zoom’s speech recognition feature to transcribe the audio. To enable the transcription, we had to turn on the audio recording. Participants were informed about Zoom’s transcription and recording features, and a pop-up notification appeared when these features were activated. After each interview, the research team checked the transcription using the recording, then immediately and permanently deleted the recording. As such, we did not store any audio recordings. In this process, we have removed all personal information (e.g., names) from the transcription, ensuring that none of the transcriptions are personally identifiable. Participants provided informed consent before the study and retained the right to withdraw at any stage. Throughout all phases of data collection (Sections 3, 4, 5), participants had the right to withdraw at any time.
Acknowledgement
We thank the anonymous reviewers and the shepherd for their invaluable feedback and the participants for their kind involvement in our studies.
Appendix A 15 Privacy-Related Design Decisions
TABLE XI: We developed a codebook of 15 privacy-related design decisions to assess the privacy-related design decisions covered in the sketches and the communication effectiveness between the sketch creator and interpreter.
Procedure Design Decision Explanation Example
Data Collection Collected Personal Data Personal data collected to achieve goals SSN, transaction history, photo, IP address
Data Provider Individuals from whom data is collected User, attendee
Collection Purpose Purpose for data collection Prediction, decision making
Data Retention Stored Data Personal data being stored Transaction history, photo
Storage Approach Data storage implementation details Data stored on cloud server for 3 months
Post-Storage Action Action taken after current data storage ends Data transferred to an archive server after being deleted on the main server
Data Processing Processing Input Personal data to be processed Transaction history, photo
Processing Output Output data after processing Risk score derived from user’s history, facial features extracted from photo
Processing Approach Data processing implementation details Facial recognition algorithm deployed on server
Data Access Accessed Raw Data Collected raw data to be accessed SSN, transaction history, photo
Accessed Output Processed data to be accessed Risk score, facial features
Access Approach Data access details Meeting host views attendees’ attention scores in system-generated report
Choice & Notice Choice Options Options for the choice maker User selects “accept” to allow data use for marketing or “reject” to deny it
Choice Impacts Effects of the user’s choice “Accept” enables data scientist to use data in marketing analysis
Choice Notification System notification upon/after choice Pop-up message such as “Your permission has been saved”
Appendix B Patterns Occurred in Creator-Interpreter Communication
TABLE XII: We summarized six patterns in the creator-interpreter communication regarding each design decision. We present sample cases to illustrate each pattern. For instance, if the creator annotated the content of a design decision (e.g., Collected Personal Data) as “SSN”, while the interpreter didn’t annotate, we would code the pattern as (E) Only the creator responded.
# Pattern Creator’s Response Interpreter’s Response
A Full alignment SSN, transaction history SSN, transaction history
B Partial alignment SSN, transaction history SSN, date of birth
C No alignment SSN Transaction history
D Only the interpreter responded SSN
E Only the creator responded SSN
F Neither responded
Appendix C Definition of Interpretation Recall and Precision in Section 5.3.2
TABLE XIII: We define precision, recall, and F1 score based on the six patterns occurred in creator-interpreter communication. Specifically, the weight of partial alignment ($\epsilon$) we used in Table VIII is 0.5. We also validated our results with $\epsilon$ = 0 in Table XIV.
Term Explanation Formula
True Positive (TP) Interpreter’s response fully or partially aligned with the creator’s response (partial alignment weighted by $\epsilon$). #$\#A+\epsilon\cdot\#B$
False Positive (FP) Interpreter responded when the creator did not, including the unaligned ($1-\epsilon$) part of $B$. #$\#D+(1-\epsilon)\cdot\#B$
False Negative (FN) The creator responded, but the interpreter either did not respond or misinterpreted the response. #$\#C+\#E$
True Negative (TN) Neither the creator nor the interpreter responded. #$\#F$
Precision The proportion of correctly aligned responses out of all responses made by the interpreter. #$\frac{\#A+\epsilon\cdot\#B}{\#A+\epsilon\cdot\#B+\#D+(1-\epsilon)\cdot\#B}$
Recall The proportion of correctly aligned responses out of all responses made by the creator. #$\frac{\#A+\epsilon\cdot\#B}{\#A+\epsilon\cdot\#B+\#C+\#E}$
F1 Score The harmonic mean of precision and recall, balancing the two. $2\cdot\frac{\textit{Precision}\cdot\textit{Recall}}{\textit{Precision}+\textit{Recall}}$
TABLE XIV: In Table VIII, we used a partial alignment weight of $\epsilon=0.5$ to calculate interpretation precision, recall, and F1. To prevent potential bias, we repeated the analysis with $\epsilon=0$ (i.e., treating “partial alignment” as “no alignment”). This table shows that the differences between the two conditions remained consistent with our previous results, indicating our findings in Section 5.3.2 are valid. For each metric, we highlight the higher value and indicate the statistical significance between the two conditions using the Wilcoxon Signed-Rank tests (${}^{*}p~{}\textless~{}0.05$, ${}^{**}p~{}\textless~{}0.01$, ${}^{***}p~{}\textless~{}0.001$), all showing large effect sizes ($r~{}\textgreater~{}0.5$).
Design Decision Vocabulary-Guided Sketches Heuristic-Guided Sketches
Precision Recall F1 Precision Recall F1
Collected Personal Data 81.0% 89.5% 85.0% 90.9% 100.0% 95.2%
Data Provider 59.1% 61.9% 60.5% 87.0% 90.9% 88.9%
Collection Purpose 22.2% 23.5% 22.9% 50.0% 58.8% 54.1%
Stored Data 18.8% 17.6% 18.2% 77.8% 77.8% 77.8%
Storage Approach 27.8% 35.7% 31.3% 68.4% 76.5% 72.2%
Post-Storage Action 6.7% 14.3% 9.1% 37.5% 42.9% 40.0%
Processing Input 23.5% 22.2% 22.9% 55.0% 61.1% 57.9%
Processing Output 42.1% 44.4% 43.2% 82.4% 93.3% 87.5%
Processing Approach 16.7% 18.2% 17.4% 53.3% 57.1% 55.2%
Accessed Raw Data 14.3% 16.7% 15.4% 54.5% 50.0% 52.2%
Accessed Output 41.7% 35.7% 38.5% 38.9% 50.0% 43.8%
Access Approach 27.3% 23.1% 25.0% 53.3% 50.0% 51.6%
Choice Options 36.4% 36.4% 36.4% 57.1% 66.7% 61.5%
Choice Impacts 0.0% 0.0% 0.0% 33.3% 57.1% 42.1%
Choice Notification 16.7% 25.0% 20.0% 66.7% 66.7% 66.7%
Mean 28.9% 30.9% 31.8% 60.4% 66.6% 63.1%
$p$-value ∗∗∗ ∗∗∗ ∗∗∗ ∗∗∗ ∗∗∗ ∗∗∗

### v2

@@
 Our user studies with 24 data science students revealed that simply presenting three heuristics to the participants at the beginning of the study can enhance the coverage of privacy-related design decisions in sketches, reduce the mental effort required for creating sketches, and improve the readability of the final sketches.
-00footnotetext:  00footnotetext: Equal contributions.00footnotetext: This report is an extended version of a IEEE S&P 2025 paper.
+00footnotetext:  00footnotetext: Equal contributions.00footnotetext: This report is an extended version of an IEEE S&P 2025 paper [References].
 We include Table XIV to further prove the improved communication efficiency of heuristic-guided privacy design sketches (Section 5.3.3).
@@
 Interestingly, the most lightweight method, which simply presents participants with a table of three heuristics (Table I), proved highly effective in improving the quality of novice students’ sketches despite its simplicity.
-We conducted a detailed experiment to validate the effectiveness of the heuristic-based approach (Section 5).
+We conducted a between-subjects experiment to validate the effectiveness of the heuristic-based approach (Section 5).
 We recruited a total of 24 participants, with each participant creating a few sketches and interpreting the sketches from other participants.
@@
 2.2 Education for Privacy Design
-Current privacy design education primarily focuses on privacy concepts [16] and design methodologies [17, 18], with most educators using case studies [17, 19] as the main form of instruction.
+Current privacy design education primarily focuses on privacy concepts [16] and design methodologies [17, 18, 19], with most educators using case studies [17, 20] as the main form of instruction.
 For instance, some classes include a privacy review procedure [21, 20], where students assess smartphone privacy notices, identify issues, and propose design improvements in brief essays.
@@
 Additionally, educators have expanded the instruction of privacy design to a broader audience, including industry professionals and policymakers [24].
-In contrast, our work aims to complement existing curricula by exploring methods to teach data science students how to sketch privacy designs, equipping them with privacy design communication skills.
+In contrast, our work aims to complement existing curricula by exploring methods to teach students how to sketch privacy designs, equipping them with privacy design communication skills.
 TABLE II: Task scenarios used in our study.
@@
 2.3 Diagrams for Privacy Design
-Practitioners often use a combination of diagrams to communicate privacy considerations in system designs [3, 4, 43].
+Practitioners often use a combination of diagrams to communicate privacy considerations in system designs [3, 4, 44, 45].
 Research has shown that these diagrams help developers visualize system architecture by simplifying complex design elements, making it easier to communicate and design privacy-aware applications [4].
@@
 While existing studies shed light on the use of diagrams for communicating privacy considerations in system design, their focus has been on professional developers.
-However, how to teach these diagramming techniques to novice data science students remains unclear.
+How to teach these diagramming techniques to novice data science students remains unclear.
 Our work aims to address this gap by using sketches as the lens to investigate the specific challenges students face when communicating privacy design through diagrams and provide insights on targeted guidance to address these difficulties.
@@
 Prior research has explored many approaches to support experienced developers in privacy design [48, 49].
-For instance, some studies have offered insights into how to help developers make privacy design choices by identifying the common challenges in understanding and implementing privacy considerations [48, 49].
+For instance, some studies have offered insights into how to help developers make privacy design choices by identifying the common challenges in understanding and implementing privacy considerations [50, 51, 52].
 Additionally, various tools have been developed to support developers by embedding privacy practices directly into development workflows.
@@
 Instead of describing each DataAction in natural language, students could select predefined attribute options.
-For example, when specifying data source to Collect, students could choose ‘direct’ for newly generated data or ‘secondary’ for pre-existing data.
+For example, when specifying data source to Collect, students could choose “direct” for newly generated data or “secondary” for pre-existing data.
 4.1.2 Tool Support
@@
 All participants completed the warm-up within 20 minutes (Mean = 18.5 minutes, SD = 1.2 minutes) and reached satisfaction with each design within 15 minutes (Mean = 13.6 minutes, SD = 2.9 minutes).
-First, we observed a significant reduction in completion time with the vocabulary-based approach (14.3 v.s.
+First, we observed a significant reduction in completion time with the vocabulary-based approach (14.3 vs.
 18.6 minutes in Section 4.2.4), as participants reported that the warm-up session helped them become familiar with the method.
@@
 Study Procedure and Apparatus.
-We followed a between-subject study design.
+We followed a between-subjects study design.
 Participants received either vocabulary-based or heuristic-based guidance for their tasks.
@@
 These time limits were decided based on the task completion time in the last experiment (Section 4.3.3).
-Then, each participant was asked to interpret two sketches created by previous participants (Table 5.2.4).
+Then, each participant was asked to interpret two sketches created by previous participants (Table 5.3).
 Finally, we interviewed participants about what hindered or assisted them in their creation and interpretation tasks.
@@
 Comparing high-dimensional sketches is challenging.
-To address this, we developed a codebook (Appendix Table XI) to annotate privacy-related design decisions covered in the sketches.
+To address this, we developed a codebook (Appendix Table LABEL:tab:design_hints) to annotate privacy-related design decisions covered in the sketches.
 We then used the annotated decisions to assess the design coverage in sketch creation and the communication effectiveness between the sketch creators and interpreters (Sections 5.2, 5.3).
@@
 We recorded their completion time and asked them to fill out a NASA TLX survey [72] after each creation.
-We presented participants with a list of privacy-related design decisions (Table XI) and asked them to describe each decision they had included, annotating the corresponding content on their sketches.
+We presented participants with a list of privacy-related design decisions (Table LABEL:tab:design_hints) and asked them to describe each decision they had included, annotating the corresponding content on their sketches.
 We didn’t set time limits for this process.
@@
 There is no perfect sketch for the creation task, as participants may have different designs for the same scenario, which may evolve during sketching.
-So, we use the codebook (Table XI) to systematically capture the unique design decisions reflected in each sketch.
+So, we use the codebook (Table LABEL:tab:design_hints) to systematically capture the unique design decisions reflected in each sketch.
 To assess the quality of the sketches, we documented the privacy design decisions annotated by participants, counted the total number of decisions included in each sketch, and calculated the coverage percentage of each design decision across all sketches in each condition.
-Next, we conducted a thematic analysis [70] of post-task interview transcripts and notes we took during study.
-First, two researchers independently coded eight transcripts (four from heuristic group and four from vocabulary group).
+Next, we conducted a thematic analysis [73] of post-task interview transcripts and notes we took during the study.
+First, two researchers independently coded eight transcripts (four from the heuristic group and four from the vocabulary group).
 After discussing and incorporating codes, we created a codebook that the two researchers agreed on.
@@
 5.2.3 Quantitative Results
+TABLE VII: In the second creation task, heuristic-guided participants reported more positive responses than they did in the first task, a trend not observed in the vocabulary group.
+We present NASA TLX results (scale of 1 to 5) as “median (mean $\pm$ standard deviation),” and highlight second task responses if they are more positive than the first.
+“” indicates statistical significance ($p~\textless~0.05$) under Wilcoxon Signed-Rank test.
+$\downarrow$” denotes that a lower value is a more positive outcome.
+Condition Task Order Mental Demand $\downarrow$ Physical Demand $\downarrow$ Temporal Demand $\downarrow$ Performance $\uparrow$ Effort $\downarrow$ Frustration $\downarrow$
+Vocabulary First 3.5 (3.17 $\pm$ 1.03) 4.0 (3.65 $\pm$ 1.23) 4.0 (3.33 $\pm$ 0.98) 3.0 (3.00 $\pm$ 0.74) 4.0 (3.83 $\pm$ 0.39) 2.0 (2.50 $\pm$ 0.90)
+Second 3.5 (3.33 $\pm$ 0.78) 4.0 (3.81 $\pm$ 1.31) 3.5 (3.17 $\pm$ 0.94) 3.0 (2.92 $\pm$ 1.08) 3.5 (3.50 $\pm$ 0.80) 3.0 (2.67 $\pm$ 0.89)
+Heuristic First 4.0 (3.67 $\pm$ 0.89) 3.5 (3.48 $\pm$ 0.97) 4.0 (3.58 $\pm$ 0.79) 3.0 (2.92 $\pm$ 0.67) 3.5 (3.58 $\pm$ 0.67) 3.0 (3.00 $\pm$ 1.04)
+Second 2.5 (2.92 $\pm$ 1.24) 3.0 (2.97 $\pm$ 0.82) 3.0 (3.00 $\pm$ 1.41) 3.5 (3.25 $\pm$ 0.87) 3.0 (3.17 $\pm$ 1.19) 2.0 (2.33 $\pm$ 1.07)
 We found that heuristics could help students sketch privacy designs more effectively. Their designs also demonstrated higher quality.
 Improved Sketching Efficiency.
-As shown in Table VI, heuristic-guided participants reached their satisfied designs faster (average of 11.53 v.s.
-14.80 minutes) and covered more privacy-related design decisions (average of 11.5 v.s.
+As shown in Table VI, heuristic-guided participants reached their satisfied designs faster (average of 11.53 vs.
+14.80 minutes) and covered more privacy-related design decisions (average of 11.5 vs.
 8.9 per sketch), compared to the vocabulary group.
@@
 This difference was statistically significant ($p~\textless~0.01{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},r=0.69}$) under the Wilcoxon Signed-Rank test [77].
-TABLE VII: In the second creation task, heuristic-guided participants reported more positive responses than they did in the first task, a trend not observed in the vocabulary group.
-We present NASA TLX results (scale of 1 to 5) as “median (mean $\pm$ standard deviation),” and highlight second task responses if they are more positive than the first.
-“” indicates statistical significance ($p~{}\textless~{}0.05$) under Wilcoxon Signed-Rank test.
-$\downarrow$” denotes that a lower value is a more positive outcome.
-Condition Task Order Mental Demand $\downarrow$ Physical Demand $\downarrow$ Temporal Demand $\downarrow$ Performance $\uparrow$ Effort $\downarrow$ Frustration $\downarrow$
-Vocabulary First 3.5 (3.17 $\pm$ 1.03) 4.0 (3.65 $\pm$ 1.23) 4.0 (3.33 $\pm$ 0.98) 3.0 (3.00 $\pm$ 0.74) 4.0 (3.83 $\pm$ 0.39) 2.0 (2.50 $\pm$ 0.90)
-Second 3.5 (3.33 $\pm$ 0.78) 4.0 (3.81 $\pm$ 1.31) 3.5 (3.17 $\pm$ 0.94) 3.0 (2.92 $\pm$ 1.08) 3.5 (3.50 $\pm$ 0.80) 3.0 (2.67 $\pm$ 0.89)
-Heuristic First 4.0 (3.67 $\pm$ 0.89) 3.5 (3.48 $\pm$ 0.97) 4.0 (3.58 $\pm$ 0.79) 3.0 (2.92 $\pm$ 0.67) 3.5 (3.58 $\pm$ 0.67) 3.0 (3.00 $\pm$ 1.04)
-Second 2.5 (2.92 $\pm$ 1.24) 3.0 (2.97 $\pm$ 0.82) 3.0 (3.00 $\pm$ 1.41) 3.5 (3.25 $\pm$ 0.87) 3.0 (3.17 $\pm$ 1.19) 2.0 (2.33 $\pm$ 1.07)
 5.2.4 Qualitative Findings
@@
 Mean 59.2% 61.8% 37.4% 39.7% 38.2% 76.7% 64.4% 69.0% 76.3% 72.2%
-[…]
+$p$-value ∗∗ $(n.s.)$ ∗∗∗ ∗∗∗ ∗∗∗ ∗∗ $(n.s.)$ ∗∗∗ ∗∗∗ ∗∗∗
+5.3 Interpretation Tasks
 TABLE IX: Interpretation task schedule.
@@
 P16 P17 P18 P23 P22 P24
-TABLE X: Participants perceived significantly lower workloads when interpreting heuristic-guided sketches than vocabulary-guided ones.
-We present their responses to NASA TLX survey as median (mean $\pm$ standard deviation), with results for heuristic-guided sketches highlighted.
-We annotate statistically significant improvements based on the Wilcoxon Signed-Rank tests (${}^{*}p~{}\textless~{}0.05$, ${}^{**}p~{}\textless~{}0.01$, ${}^{***}p~{}\textless~{}0.001$), all with medium to near large effect sizes ($r:0.3\sim 0.5$).
-$\downarrow$” indicates that a lower value is more positive.
-Condition Mental Demand $\downarrow$ Physical Demand $\downarrow$ Temporal Demand $\downarrow$ Performance $\uparrow$ Effort $\downarrow$ Frustration $\downarrow$
-Vocabulary 3.0 (3.41 $\pm$ 1.10)∗∗∗ 3.5 (3.66 $\pm$ 0.92)∗∗ 3.0 (2.82 $\pm$ 1.18) 3.0 (2.91 $\pm$ 1.15)∗∗ 3.5 (3.32 $\pm$ 1.04)∗∗∗ 3.0 (3.23 $\pm$ 1.07)∗∗∗
-Heuristic 2.0 (2.29 $\pm$ 0.86)∗∗∗ 2.5 (2.61 $\pm$ 0.88)∗∗ 2.0 (2.17 $\pm$ 0.87) 4.0 (3.75 $\pm$ 0.79)∗∗ 2.0 (2.25 $\pm$ 0.85)∗∗∗ 2.0 (2.04 $\pm$ 0.81)∗∗∗
-5.3 Interpretation Tasks
 5.3.1 Data Collection
 As shown in Figure 6, each participant completed two interpretation tasks, each with a 7-minute time limit.
-In each task, they received a randomized list of privacy-related design decisions (excluding the “Procedure” column in Table XI) and an unannotated copy of a sketch created by a prior participant.
+In each task, they received a randomized list of privacy-related design decisions (excluding the “Procedure” column in Table LABEL:tab:design_hints) and an unannotated copy of a sketch created by a prior participant.
 We then asked them to annotate and orally describe the privacy-related design decisions they identified.
 This annotation is similar to what they have done during the creation tasks (Section 5.2.1).
-Table 5.2.4 outlines the interpretation task order, where each participant interpreted two sketches of different scenarios: one created by a heuristic-guided participant and another from the vocabulary group.
+Table 5.3 outlines the interpretation task order, where each participant interpreted two sketches of different scenarios: one created by a heuristic-guided participant and another from the vocabulary group.
 We maintained scenario order consistent with the creation tasks (Table 5.2.1) to avoid biases from the most recent scenario they had sketched.
@@
 The qualitative analysis shared the same procedure as described in Section 5.2.2, and we will report the findings of interpretation tasks in Section 5.3.4.
+TABLE X: Participants perceived significantly lower workloads when interpreting heuristic-guided sketches than vocabulary-guided ones.
+We present their responses to NASA TLX survey as median (mean $\pm$ standard deviation), with results for heuristic-guided sketches highlighted.
+We annotate statistically significant improvements based on the Wilcoxon Signed-Rank tests ($^{*}p~\textless~0.05$, $^{**}p~\textless~0.01$, $^{***}p~\textless~0.001$), all with medium to near large effect sizes ($r:0.3\sim 0.5$).
+$\downarrow$” indicates that a lower value is more positive.
+Condition Mental Demand $\downarrow$ Physical Demand $\downarrow$ Temporal Demand $\downarrow$ Performance $\uparrow$ Effort $\downarrow$ Frustration $\downarrow$
+Vocabulary 3.0 (3.41 $\pm$ 1.10)∗∗∗ 3.5 (3.66 $\pm$ 0.92)∗∗ 3.0 (2.82 $\pm$ 1.18) 3.0 (2.91 $\pm$ 1.15)∗∗ 3.5 (3.32 $\pm$ 1.04)∗∗∗ 3.0 (3.23 $\pm$ 1.07)∗∗∗
+Heuristic 2.0 (2.29 $\pm$ 0.86)∗∗∗ 2.5 (2.61 $\pm$ 0.88)∗∗ 2.0 (2.17 $\pm$ 0.87) 4.0 (3.75 $\pm$ 0.79)∗∗ 2.0 (2.25 $\pm$ 0.85)∗∗∗ 2.0 (2.04 $\pm$ 0.81)∗∗∗
 5.3.3 Quantitative Results
@@
 All participants finished their interpretation within the time limit.
-There was no significant difference in interpretation time (4.9 minutes for heuristic-guided sketches v.s.
+There was no significant difference in interpretation time (4.9 minutes for heuristic-guided sketches vs.
 5.1 minutes for vocabulary-guided sketches).
@@
 First, heuristic-guided sketches achieved higher interpretation performance across metrics, with an increase of 31.6% in precision (i.e., increased fraction of correct interpretations) and 36.6% in recall (i.e., increased fraction of creator’s design decisions that were accurately interpreted).
-At the level of a complete privacy design, heuristic-guided sketches demonstrated higher per-sketch interpretation performance, again with a higher precision (79.1% v.s.
-44.6% for the vocabulary group), reflecting the proportion of matches among all interpretations in a sketch, and higher recall (71.1% v.s.
+At the level of a complete privacy design, heuristic-guided sketches demonstrated higher per-sketch interpretation performance, again with a higher precision (79.1% vs.
+44.6% for the vocabulary group), reflecting the proportion of matches among all interpretations in a sketch, and higher recall (71.1% vs.
 40.5%), representing the percentage of alignments among the creator’s all design decisions.
@@
 Furthermore, the Multi-layer Representation heuristic helps students manage their sketches from a macro perspective, allowing them to approach designs strategically.
-Sketches v.s.
+Sketches vs.
 Diagrams.
@@
 Our current study focuses on facilitating the communication of privacy design concepts through sketches.
-Future research may further explore the benefits of sketches in other security and privacy applications.
+Future research may explore the benefits of sketches in other security and privacy applications.
 For example, participants may develop a stronger awareness of privacy risks during sketching [84, 85, 86, 87].
@@
 The creation task has no ground truth.
-While our codebook (Table XI) provided a structured approach for labeling and interpreting data, it may have constrained the granularity of our analysis.
+While our codebook (Table LABEL:tab:design_hints) provided a structured approach for labeling and interpreting data, it may have constrained the granularity of our analysis.
 By relying on predefined categories, we may have overlooked subtle contextual differences or nuances in participant responses.
@@
 We thank the anonymous reviewers and the shepherd for their invaluable feedback and the participants for their kind involvement in our studies.
+This research is in part supported by a gift from Solix Technologies, the National Science Foundation CNS-2341187, CNS-2426397, a Google PSS Faculty Award, and a Meta Research Award.
 Appendix A 15 Privacy-Related Design Decisions
@@
 Term Explanation Formula
-True Positive (TP) Interpreter’s response fully or partially aligned with the creator’s response (partial alignment weighted by $\epsilon$). #$\#A+\epsilon\cdot\#B$
-False Positive (FP) Interpreter responded when the creator did not, including the unaligned ($1-\epsilon$) part of $B$. #$\#D+(1-\epsilon)\cdot\#B$
-False Negative (FN) The creator responded, but the interpreter either did not respond or misinterpreted the response. #$\#C+\#E$
-True Negative (TN) Neither the creator nor the interpreter responded. #$\#F$
-Precision The proportion of correctly aligned responses out of all responses made by the interpreter. #$\frac{\#A+\epsilon\cdot\#B}{\#A+\epsilon\cdot\#B+\#D+(1-\epsilon)\cdot\#B}$
-Recall The proportion of correctly aligned responses out of all responses made by the creator. #$\frac{\#A+\epsilon\cdot\#B}{\#A+\epsilon\cdot\#B+\#C+\#E}$
+True Positive (TP) Interpreter’s response fully or partially aligned with the creator’s response (partial alignment weighted by $\epsilon$). $A+\epsilon\cdot\#B$
+False Positive (FP) Interpreter responded when the creator did not, including the unaligned ($1-\epsilon$) part of $B$. $D+(1-\epsilon)\cdot\#B$
+False Negative (FN) The creator responded, but the interpreter either did not respond or misinterpreted the response. #C+#E\#C+\#E
+True Negative (TN) Neither the creator nor the interpreter responded. #F\#F
+Precision The proportion of correctly aligned responses out of all responses made by the interpreter. $\frac{\#A+\epsilon\cdot\#B}{\#A+\epsilon\cdot\#B+\#D+(1-\epsilon)\cdot\#B}$
+Recall The proportion of correctly aligned responses out of all responses made by the creator. $\frac{\#A+\epsilon\cdot\#B}{\#A+\epsilon\cdot\#B+\#C+\#E}$
 F1 Score The harmonic mean of precision and recall, balancing the two. $2\cdot\frac{\textit{Precision}\cdot\textit{Recall}}{\textit{Precision}+\textit{Recall}}$
@@
 $p$-value ∗∗∗ ∗∗∗ ∗∗∗ ∗∗∗ ∗∗∗ ∗∗∗
+Appendix D Meta-Review
+The following meta-review was prepared by the program committee for the 2025 IEEE Symposium on Security and Privacy (S&P) as part of the review process as detailed in the call for papers.
+D.1 Summary
+The paper presents a user study of n=24 data science students (both undergraduate and graduate) who were asked to sketch privacy designs of a set of privacy scenarios after being given three heuristics to follow at the beginning of the study.
+The paper describes two pre-studies.
+The first pre-study was a need-finding study (n=12) that found that students lacked the vocabulary needed to sketch privacy designs effectively, the sketching may require multiple iterations, and the students had difficulty planning the sketch space in advance.
+The second pre-study is a teaching methods experiment that investigates 3 methods: object-oriented diagramming, vocabulary-based sketching, and heuristic-based sketching.
+The findings suggested that heuristic-based and vocabulary-based sketching would be the best for data science students, and those methods are examined in the main study.
+The main study (n=24) compared the two methods in a sketching task, followed by a post-task interview.
+The findings suggest that the simple heuristic method performed best in sketching efficiency and broader coverage of privacy-related design decisions and had a reduced learning curve.
+A heuristic-based approach combined with the multi-layer representation is recommended for privacy design sketches conducted by data science students.
+D.2 Scientific Contributions
+  • Provides a Valuable Step Forward in an Established Field
+  • Establishes a New Research Direction
+D.3 Reasons for Acceptance
+  1. 1.
+    Provides a Valuable Step Forward in an Established Field.
+The paper contributes to the research on sketching privacy designs and offers insights into how we should teach data science students about data privacy design.
+  2. 2.
+    Establishes a New Research Direction.
+This paper presents the design and evaluation of a novel approach to prompting data science students to consider privacy in their design.
+The proposed heuristic-based prompting offers a novel direction of investigation for future work.